The half-life of 14C is 5730 yr. Find the number of disintegrations emitted by 1.0 mol of 14C in 1 minute. (Note: 1 year = 365.24 day). Express your answer using two significant figures.
Answer #1

activity = KN C-14 N = 6.02x1023 atoms ty2 = 5730 yr = 5730 X 365.24 x 24 x 60 = 3.01x109 min K = In2/t1/2 = 0.693 / 3.01x109